OpenStudy (anonymous):

Solve x^4 = 36x^2. Identify any double roots. A. {0 (double root), 6, –6} B. {–3, 3, 2 (double root)} C. {0, 3, 2} D. {6, –6}

OpenStudy (anonymous):

\[ x^4=36x^2\implies x^4-36x^2=0 \implies x^2(x^2-36)=0\implies x^2(x-6)(x+6)=0 \]

OpenStudy (anonymous):

What would be the right answer? :(